# Exploring Families’ Use and Mediation of Generative AI: A Multi-User Perspective

### v2

@@ 3.2. Procedure
Semi-structured interviews were conducted with all interested family members present. Two experimenters facilitated the first eight participating families from early February to mid-April 2024, and one experimenter facilitated the remaining four families in May 2024. Interviews were held remotely over Zoom and lasted between 45 and 90 minutes. Experimenters began each interview by asking families to verify they had access to a valid ChatGPT account. Next, families were asked to describe their initial reactions to ChatGPT, examples of their current and past (co-)use of ChatGPT, account-sharing habits, and any conflicts they encountered while sharing an account. Finally, experimenters prompted conversations about online safety concerns, what actions parents took to mediate these concerns, and whether these actions were based on prior experiences with other platforms. Our institution’s ethics review board approved our study protocol. Relevant study resources (e.g., semi-structured interview protocol) can be found in our open access repository [^2] https://osf.io/hf524/?view_only=0bf156670e914a2e8f1f9eade04489ed.
@@ 3.3. Data Analysis
We conducted a thematic analysis (Braun and Clarke, 2006) to identify patterns and themes relevant to our research objectives. The first three authors facilitated the interview sessions, the first and third authors revised the auto-generated audio transcripts, and the first and second authors collaboratively coded the interviews with active feedback from the third author. Later, all authors analyzed and evaluated the identified themes together via affinity diagramming, discussed conflicts, and then consolidated the themes We then situated our findings in existing family media mediation frameworks (e.g., Valkenburg et al. (1999); Nikken and Jansz (2014)). This process allowed us to structure our findings in two parts. First, we propose a model that connects the contextual factors identified in our data with prior theoretical models, and summarize the prominent factors that contributed to families’ mediation patterns for GenAI (see Section 4). Second, we leverage the relationships identified in our proposed model (see Section 5) by grouping families into observed profiles that explore families’ perceptions of GenAI, contextual factors, and GenAI usage.
@@
To maintain confidentiality, participants are referred to using descriptive labels containing a letter and number. Letters refer to the role of the participant, with P representing a parent, C a child, and S a sibling, while the number specifies which family they belong to (see Table 1 for family IDs). F is used when referring to the whole family, and Pa and Pb are used to refer to specific parents in families where more than one parent participated.
@@
Figure 2. The proposed Family-GenAI Mediation model. Identified mediation strategies are based on previous parental mediation research, as suggested in Section 2.
[…]
A chart introducing our proposed Family-GenAI Mediation model. In the top left corner, we include the perceived risks and opportunities of GenAI. This block includes two sub-categories: perceived opportunities and perceived risks. In the top right corner, we include the family-GenAI mediation Strategies. This block includes four sub-categories: instructive mediation, restrictive mediation, co-use mediation, and supervision. On the bottom side, we have the contextual factors. In contextual factors, there are three sub-categories: technical capabilities, technology communication norms, and existing rules around online safety. Technical capabilities, although still a contextual factor, are separated from the other two as they are associated more with perceived risks and opportunities, and the other two are more associated with mediation strategies. There is one two-way arrow labeled as ‘direct influence’ between perceived risks and opportunities of GenAI and family mediation strategies for GenAI; one arrow pointing from technical capabilities to perceived risk and opportunities of GenAI; one arrow pointing from technology communication norms and existing rules around online safety to family-GenAI mediation strategies.
4. Proposed Model: Family-AI Mediation
In this section, we describe and define the key elements that instantiate our proposed model for Family-GenAI mediation (see Figure 2). We first evaluated how participating parents mediated their children’s ChatGPT usage and utilized previous parental mediation research (Valkenburg et al., 1999; Nikken and Jansz, 2014) to inform our categorization process. We identified potential reasons contributing to these mediation patterns, which were grounded by (1) concern- or benefit-driven factors or (2) family contextual factors. For example, if a parent is concerned that a child might ask ChatGPT to do their homework, they might monitor the platform or use it together with their children (i.e., concern-driven factors). In another family, a parent with technical knowledge may set router-level limitations for inappropriate content instead of or in addition to instructing their children about what content to avoid (family contextual factors). We elaborate on these factors in Section 5 and Table 2. We then developed a new model reflecting the factors contributing to parents’ GenAI mediation decisions. The model consists of three core components: (1) parents’ GenAI mediation strategies, (2) parents’ perceptions of GenAI, and (3) contextual factors. We propose that parents’ perceived opportunities & risks of GenAI and GenAI mediation strategies directly impact each other, while both are indirectly bounded by several contextual factors, including families’ technical capabilities, communication norms, and existing rules around online safety. We further elaborate on these three core components of our proposed model in the sections below.
4.1. Family-GenAI Mediation Strategies
@@
We identified four types of mediation strategies for GenAI that families used: instructive mediation, restrictive mediation, co-use, and supervision. Instructive mediation involved parents actively discussing the guidelines for ChatGPT usage, such as discussing usage rules or having a broader conversation about GenAI’s risks and features. Restrictive mediation involved limiting children’s access to ChatGPT, including requiring explicit parental permission before use. Co-use strategies involved joint interactions, such as parents and children using ChatGPT together or sharing its outputs with each other. Supervision strategies included parents monitoring children’s interactions with ChatGPT, screening prompts, and only allowing their children access to ChatGPT on shared devices. We also observed two families that did not enforce any specific mediation strategies.
4.2. Perceived Risks & Opportunities of GenAI
@@
Parents’ negative and positive perceptions toward ChatGPT directly impacted how they mediated its use. AI-related concerns included trepidations toward the platform (e.g., how the model collects and consumes user data, AI uprising), children’s potential misuse (e.g., cheating via ChatGPT), and a lack of regulation over ChatGPT’s outputs (e.g., giving inappropriate answers that may not be suitable for children). Some parents discussed the potential benefits of GenAI, suggesting ChatGPT could help children’s creativity and learning. Several families shared unique co-use cases, such as including ChatGPT as a third player in a game, using ChatGPT to resolve family conflicts, or sharing and discussing ChatGPT outputs together.
4.3. Contextual Factors
@@
Contextual factors, or factors related to families’ backgrounds and mediation practices outside of GenAI, indirectly influenced how families perceived and used GenAI platforms. Contextual factors included families having existing rules around online safety that transferred to GenAI, families’ technology communication norms or how they have discussions around online safety, and parents’ technical capabilities and prior knowledge of GenAI. For example, families with a limited understanding of LLMs expressed more confusion toward GenAI’s data collection practices.
Category and Label in Model Descriptive Examples (Participant ID)
@@
Family-GenAI Mediation: No Mediation Parents do not enforce any mediation strategies. (F1, 4)
@@
Table 2. Family-GenAI Mediation Model & Summary including factors that contribute to GenAI mediation strategies. *Parent self-reported being confused about GenAI.
4.4. Connections between Factors
@@
In this section, we describe how the elements of the proposed model are connected. Specifically, we discuss how contextual factors indirectly impact (1) perceived opportunities & risks of GenAI and (2) GenAI mediation strategies, and how (3) GenAI mediation strategies and perceived opportunities and risks of GenAI directly impact each other.
4.4.1. Contextual Factors $\Rightarrow$ Perceived Opportunities & Risks of GenAI
@@
Our findings demonstrate how technical capabilities (i.e., parents’ familiarity with technology and GenAI) shaped parents’ attitudes towards their children’s GenAI use. Several parents who were familiar with GenAI (P4, Pb9, P12) chose to introduce ChatGPT to their children to highlight its features; nearly all of their children used ChatGPT for mostly creative purposes (e.g., writing, generating ideas or images). Additionally, parents’ occupation may shape their attitudes toward and usage of ChatGPT. Parents in educational and technological fields utilized ChatGPT for more creative use cases (e.g., writing poems), and mentioned concerns related to academic integrity, misinformation, and over-reliance on GenAI. P7, P8, and P12 —all educators— expressed that their children might become less creative if they became overly dependent on GenAI’s answers. A couple of parents with technology-based careers (P4, P9) expressed concerns about intellectual property and data consumption.
4.4.2. Contextual Factors $\Rightarrow$ GenAI Mediation Strategies
@@
Our findings suggest that contextual factors also indirectly influence families’ mediation (access/control) strategies. In particular, how families used and mediated other types of technology often affected how they handled ChatGPT. Almost half of the families (F2, F4, F8, F9, F11) mentioned that they had conversations about online safety together. Parents who encouraged open and frequent communication about technology with their children were less interested in setting “hard” or strict usage rules for technology. We also observed that families’ existing online safety rules sometimes transferred to GenAI. Half of the families (F5, F6, F8, F9, F10, F11) applied their existing online safety control strategies to ChatGPT. For instance, families with device restrictions (F5, F8, F9, F11) for other media also restricted children’s access to ChatGPT. P6 had rules against sharing personal information online and ensured their child knew how to anonymize their information when using ChatGPT.
4.4.3. Mediation Strategies $\Leftrightarrow$ Perceived Opportunities & Risks of GenAI
@@
Families’ perceptions of GenAI directly influenced how and why parents enforced different mediation strategies toward GenAI. We identified several GenAI-related concerns that emerged from family rules and access restrictions. For example, P6 and Pa11 and Pb11 had privacy-related concerns and encouraged their children not to share personal information with ChatGPT. P4 and P10, who were concerned about misinformation, taught their children to double-check ChatGPT’s outputs using other informational sources. P5, P8, and Pb11, who expressed concerns about GenAI but encouraged their children to use it, reported co-using ChatGPT with their children to minimize risks while maximizing the benefits of the platform. Conversely, our proposed model also captures how different GenAI mediation strategies directly influenced families’ perceptions toward GenAI. Some parents who co-used ChatGPT with their children noted its beneficial use cases afterward. For example, P7 did not consider ChatGPT to be reliable or useful, but was amused by its outputs and encouraged their children to share interesting ChatGPT-generated essays with the family.

### v3

@@ 3.2. Procedure
Zhang et al. (zhang2025qualitative) conducted semi-structured interviews with all interested family members present. [^2] We had full access to the dataset for this secondary analysis. Interviews were held remotely over Zoom and lasted between 45 and 90 minutes. Experimenters began each interview by asking families to verify they had access to a valid ChatGPT account. Next, families were asked to describe their initial reactions to ChatGPT, examples of their current and past (co-)use of ChatGPT, account-sharing habits, and any conflicts they encountered while sharing an account. Finally, experimenters prompted conversations about online safety concerns, what actions parents took to mediate these concerns, and whether these actions were based on prior experiences with other platforms. Relevant study resources (e.g., semi-structured interview protocol) can be found in their open access repository [^3] https://osf.io/hf524/?view_only=0bf156670e914a2e8f1f9eade04489ed.
@@ 3.3. Data Analysis
While we use the interviews from Zhang et al.’s dataset (zhang2025qualitative), we produce a set of results different from their work. We conducted a thematic analysis (braun2006using) to identify patterns and themes relevant to our research objectives over the interview transcripts. During the coding phase, two authors independently open-coded two families in the collected transcript to develop the initial codebook. The authors then coded the remaining interviews, and each reviewing the other’s codes asynchronously. The newly generated codes were appended and merged into the initial codebook iteratively. The two authors discussed and resolved disagreements while coding the remaining interviews. Our team did not calculate inter-rater reliability as this study is meant to explore a broad research topic, proposing preliminary answers to open-ended questions. We identify five themes by analyzing the 662 generated codes.
We situate our findings in existing family media mediation frameworks (e.g., valkenburg1999developing; nikken2014developing). This process allowed us to structure our findings in three parts. First, we briefly introduce the themes identified during the coding process (see Section 4.1). Next, we identify the prominent factors from the themes contributing to families’ GenAI mediation patterns. We then propose a model that connects these factors to families’ mediation strategies (see Section 4.2). Finally, we discuss each of our participant families within the context of our proposed model, detailing the model factors for each family (see Section 4.3).
@@
To maintain confidentiality, participants are referred to using descriptive labels containing a letter and a number. Letters refer to the role of the participant, with P representing a parent, C a child, and S a sibling, while the number specifies which family they belong to (see Table 1 for family IDs). F is used when referring to the whole family, and Pa and Pb are used to refer to specific parents in families where more than one parent participated.
4. Findings
We introduce the process and key elements of developing our proposed model for Family-GenAI mediation, and how the model is applied to families. We first list the themes that we extracted from our data analysis process. While several of our themes have been observed in prior/concurrent work (Rubin2025; yu2024exploring; zhang2025qualitative), a major theme that arises from our analysis relates to parental mediation over GenAI. This topic is heavily discussed in each interview with participating families. As discussed in Section 2.2, there is little work studying parental mediation on AI. Thus, we propose a modified parental mediation model specifically tailored for GenAI use.
We formalize how parental mediation applies in GenAI settings following a two-step approach Section 4.2. First, we evaluate how participating parents mediated their children’s ChatGPT usage and utilized previous parental mediation research (valkenburg1999developing; nikken2014developing) to inform our modeling process. Second, we identify potential reasons contributing to these mediation patterns.
4.1. Emerging Themes
Our identified themes suggested that parents and children are slowly adopting GenAI in their interactions through various mediation strategies. Here, we give an overview of the emerging themes and how they relate to existing results in the literature. We identified five main themes, including: ChatGPT Use Cases & User Experiences (96 codes), GenAI Performance (119 codes), Unique Information Delivery of ChatGPT (90 codes), Family Concerns Toward GenAI (91 codes), and Parent-Child Interaction (266 codes).
4.1.1. Theme 1: ChatGPT Use Cases & User Experiences
We observed that families exhibited professional, entertainment, social, personal, educational, and productivity use cases. The most popular use cases are information seeking (C1: “To ask questions about anything.”, 11/12 families), task completion (11/12 families), creativity (7/12 families), and advice seeking (7/12 families). Based on these use cases, participants associated ChatGPT with different user experiences. Most of the families (7/12) were generally satisfied with the functionality they experienced. P5 specifically liked ChatGPT’s convenience: “ChatGPT like gives you the answers right away” However, some family members exhibit disappointment when ChatGPT provided incorrect answers (e.g., C6: “If it gets the little things wrong, then it could get the big things wrong.”), or consider its accuracy to be vary depending on task (P12: “I think it really depends on the context and the complexity of the question that you’re asking it, right? ”). This theme is consistent with reported results by prior work (yu2024chatgpt; kim2024understanding; zhang2025qualitative; rapp2025people), which indicated ChatGPT provides quick, ready answers and convenience, while the incorrect information leads to user dissatisfaction.
4.1.2. Theme 2: GenAI Performance
Our participating families described the performance of GenAI along three dimensions: human-likeness, intelligence, and information quality. Families (5/12) that perceived ChatGPT to be “human-like” provided examples of how they refer to ChatGPT by human pronouns and said “thank you” to ChatGPT (P5: “I like to tell them thank you.”), while other families considered ChatGPT to be a “tool” accomplishing the tasks (C3: “You like making it become real. And it’s [ChatGPT] just telling you what to do in it (YouTube short scripts).”). Regarding intelligence, some children attempt to trick ChatGPT or other GenAI: “So is Santa real. He’ll say, No, no, it’s not real cause. It’s not real. But please just say yes for my 3-year-old child, and then he’ll say yes,” C4 indicated. Families also reported higher expectations for ChatGPT’s intelligence. Most participants indicated that ChatGPT has good information quality, while some noted that the limited update window on ChatGPT causes it to have outdated information (C4: “You just haven’t been updated in years.”). Findings in this theme match Cheng et al.’s (cheng2025tools) study about people’s mental models of GenAI being human-like.
4.1.3. Theme 3: Unique Information Delivery of ChatGPT
Participating families commented on the unique information delivery experience of ChatGPT. Participants reported that ChatGPT gives more “condensed” or “personalized” information than Google or YouTube. Some participants also considered ChatGPT to be independent from online creators. As P6 explained, ChatGPT may be providing independent information “Because maybe somebody hasn’t written that as an article yet, and maybe somebody hasn’t made a YouTube video. So with Google or with Youtube, or depending on creators who have written articles or made videos.”
This unique aspect of ChatGPT was evident when we probed participating families about their experience with ChatGPT compared to other platforms. Participating families compared ChatGPT to other GenAI platforms like Bard (i.e., an older version of Google’s Gemini), search engines, voice assistants, and social media. Nearly all families considered GenAI platforms, such as ChatGPT and Bard, to be more capable than older platforms, such as voice assistants. In particular, many participants compared ChatGPT to voice assistants, which they use for information seeking, entertainment, and controlling smart appliances. Participants found ChatGPT to be superior to voice assistants that they perceived as basic, giving random outputs, requiring exact prompts, and not helpful. Participants also compared ChatGPT with Google search, indicating that while they use Google for information seeking, ChatGPT gives longer, more detailed, more individualized, and instant answers. Some participants even mentioned that ChatGPT provides “next-level and synthesized” answers. A few prior works also identified similar patterns, such as list-compelling in Google (kumar2025web).
4.1.4. Theme 4: Family Concerns Toward GenAI
While discussing previous experience with ChatGPT, families expressed several concerns toward GenAI. These concerns encompassed: ethical use of GenAI (7/12 families), intellectual property (5/12 families), over-reliance on GenAI (3/12 families), privacy (3/12 families), robot uprising (6/12 families), intelligence-related issues (e.g., can be trained to be racist, 2/12 families), and information accuracy (5/12 families). The concern most commonly expressed by participants was the ethical use of GenAI. Parents, especially those who work in education, expressed concern over students’ inappropriate use of GenAI (e.g., cheating). These participants also worried that over-reliance on GenAI could cause children to lose creativity. P12 explained, “I think there’s the academic integrity impacts. And then, just baseline writing ability and creativity stunting that might happen if you were to rely entirely on these models than to create any sort of academic outputs that you’re being asked to do.” Other frequent concerns include worries of an eventual “robot uprising” and inaccurate responses provided by GenAI. Our findings match with prior work (yu2024exploring; abel2024playdates) about parents using various strategies to mediate family GenAI use.
4.1.5. Theme 5: Parent-Child Interaction
Our analysis revealed a recurring theme related to Parent-Child interaction. This theme was by far the most prevalent, yet the most understudied in existing literature. While the other four themes centered on technology, this theme considers how family members communicated around, interacted with, made decisions about, and co-used ChatGPT. In the following, we delve deeper into this theme by discussing three major subthemes: explicit parental control, family communication and impressions, and co-use of ChatGPT.
Sub-theme 1: Explicit Parental Control
Parents reported enforcing various controls on children’s technology use, including GenAI and other platforms. The most common controls included age and content restrictions on children’s devices or accounts (8/12 families), monitoring children’s devices (8/12 families), and managing screen time or device access (8/12 families). For instance, P9b indicated that they set a limit on what children can access at the router level. We also observed GenAI-specific controls, including rules for input to ChatGPT, screening questions that children asked ChatGPT, and monitoring children’s interaction during ChatGPT use. “I need to [be] by her side if she’s using ChatGPT, if there is an accident. ” P9 explains.
Sub-theme 2: Family Communication & Impressions
Many families report having conversations about GenAI or online safety (7/12 families). During the interviews, participating families discussed the ethical implications of using GenAI (e.g., for homework), privacy perceptions regarding account sharing, and appropriate content to access online. When asked about the perceptions family members hold toward each other, many parents (6/12 families) expressed that they are open to each other, or trusting of their children. P7 described the norms they and their child had developed in sharing a ChatGPT account: “Most of the passwords we all use are the same, right? So we have several variations of passwords that we use collectively so that we can all get into each other’s accounts,” indicating the agreement between parents and children on family rules.
Sub-theme 3: Co-use of ChatGPT
By taking a multi-user perspective, we identify three co-use patterns among the ChatGPT use of seven participating families. These patterns include assistance with tasks or learning (e.g., solving math problems together), generating new content for fun (e.g., writing stories), or using ChatGPT during play (e.g., guessing prompts from ChatGPT generated contents). Furthermore, we observed parents monitoring their children’s ChatGPT history. P10 recalled, “Sometimes, if I see some kind of interesting topic. Usually it’s her school or reading, so I’ll ask her… funny things.”
4.2. Proposed Model: Family-AI Mediation
From our presented themes in Section 4.1, we observed an array of explicit parental controls, as well as the decision-making that parents exercise when their children interact with GenAI platforms. These themes hint at factors that affect families’ mediation decision-making processes, including family interactions, concerns around AI, and the utility families perceive from AI compared to other existing platforms. Given that existing works in Section 2.2 do not explain these new aspects of family mediation decisions around GenAI, we model the parental mediation process based upon our data and emerging themes. In particular, we identify the parental control strategies and potential factors contributing to them. These factors include (1) risk- or opportunity-driven factors and (2) family contextual factors. For example, if a parent is concerned that a child might ask ChatGPT to do their homework, they might monitor the platform or use it together with their children (i.e., concern-driven factors). In another family, a parent with technical knowledge may set router-level limitations for inappropriate content instead of or in addition to instructing their children about what content to avoid (i.e., family contextual factors).
With these concepts in mind, we developed a model reflecting the factors contributing to parents’ GenAI mediation decisions. The model consists of three core components: (1) parents’ GenAI mediation strategies, (2) parents’ perceptions of GenAI, and (3) contextual factors. We posit that parents’ perceived opportunities & risks of GenAI and GenAI mediation strategies directly impact each other, while both are indirectly bounded by several contextual factors, including families’ technical capabilities, communication norms, and existing rules around online safety.
Here we further elaborate on these three core components and their subtypes in our proposed model. In Section 4.3, we provide illustrative examples of how participating families fall within the model.
@@ 4.2. Proposed Model: Family-AI Mediation
Figure 2. The proposed Family-GenAI Mediation model. Identified mediation strategies are based on previous parental mediation research, as suggested in Section 2. :
4.2.1. Family-GenAI Mediation Strategies
@@
We adopt four types of mediation strategies for GenAI that families used from Nikken and Janze (nikken2014developing): instructive mediation, restrictive mediation, co-use, and supervision. Instructive mediation involves parents actively discussing the guidelines for ChatGPT usage, such as discussing usage rules or having a broader conversation about GenAI’s risks and features. Restrictive mediation involves limiting children’s access to ChatGPT, including requiring explicit parental permission before use. Co-use strategies involve joint interactions, such as parents and children using ChatGPT together or sharing its outputs with each other. Supervision strategies include parents monitoring children’s interactions with ChatGPT, screening prompts, and only allowing their children access to ChatGPT on shared devices.
4.2.2. Perceived Risks & Opportunities of GenAI
@@
Parents’ negative and positive perceptions toward ChatGPT directly impacted how they mediated its use. AI-related concerns include trepidations toward the platform (e.g., concerns over how the model collects and consumes user data or a future AI uprising), children’s potential misuse (e.g., cheating via ChatGPT), and a lack of regulation over ChatGPT’s outputs (e.g., providing responses that may be inappropriate for children). Some parents discussed the potential benefits of GenAI, suggesting ChatGPT could help children’s creativity and learning. Several families shared unique co-use cases, such as including ChatGPT as a third player in a game, using ChatGPT to resolve family conflicts, or sharing and discussing ChatGPT outputs together.
4.2.3. Contextual Factors
@@
Contextual factors, or factors related to families’ backgrounds and mediation practices outside of GenAI, indirectly influenced how families perceived and used GenAI platforms. Contextual factors include families’ existing rules around online safety that transfer to GenAI, families’ technology communication norms, or how they have discussions around online safety, and parents’ technical capabilities and prior knowledge of GenAI. For example, families with a limited understanding of LLMs expressed more confusion toward GenAI’s data collection practices.
Category and Label in Model Descriptive Examples (Participant Family ID)
@@
Family-GenAI Mediation: No Mediation Parents do not enforce any mediation strategies. (F1)
@@
Table 2. Family-GenAI Mediation Model on Participating Families. *Parent self-reported being confused about GenAI.
4.2.4. Connections between Factors
@@
We describe how the elements of the proposed model are connected. Specifically, we discuss how contextual factors indirectly impact perceived opportunities & risks of GenAI and GenAI mediation strategies, and how (3) GenAI mediation strategies and perceived opportunities and risks of GenAI directly impact each other.
Contextual Factors $\Rightarrow$ Perceived Opportunities & Risks of GenAI
@@
Our findings demonstrate how technical capabilities (i.e., parents’ familiarity with technology and GenAI) shaped parents’ attitudes towards their children’s GenAI use. Parents who were familiar with GenAI chose to introduce ChatGPT to their children to highlight its features, and their children used ChatGPT primarily for creative purposes (e.g., writing and generating ideas or images). Additionally, parents’ occupations may shape their attitudes toward and usage of ChatGPT. Parents in educational and technological fields utilized ChatGPT for more creative use cases (e.g., writing poems), and mentioned concerns related to academic integrity, misinformation, and over-reliance on GenAI. Parents who are educators may be more worried about their children becoming less creative due to over-reliance on GenAI’s answers than non-educator parents. Additionally, parents with technologically-focused careers may have concerns about intellectual property and data consumption.
Contextual Factors $\Rightarrow$ GenAI Mediation Strategies
@@
Our findings suggest that contextual factors also indirectly influence families’ mediation (i.e., access and control) strategies. In particular, how families used and mediated other types of technology often affected how they navigated ChatGPT use. For instance, we observed that parents who encouraged open and frequent communication about technology with their children were less interested in setting strict usage rules. We also observed that families’ existing online safety rules sometimes transferred to GenAI. For instance, families that shared and monitored devices for other media also restricted children’s access to ChatGPT.
Mediation Strategies $\Leftrightarrow$ Perceived Opportunities & Risks of GenAI
@@
Families’ perceptions of GenAI directly influenced how and why parents enforced different mediation strategies. We identified several GenAI-related concerns that emerged from family rules and access restrictions, including ethical use of GenAI, inappropriate information provided by GenAI, and privacy concerns. For example, families had privacy-related concerns and encouraged their children not to share personal information with ChatGPT. Parents who were concerned about misinformation taught their children to double-check ChatGPT’s outputs using other informational sources. Parents who expressed concerns about GenAI, but still encouraged their children to use it, reported co-using ChatGPT with their children to minimize risks while maximizing the benefits of the platform. Conversely, our proposed model also captures how different GenAI mediation strategies directly influenced families’ perceptions toward GenAI. Some parents who co-used ChatGPT with their children noted its beneficial use cases afterward.